RHIC Startup:
Three separate ring accesses yesterday. Blue abort kicker had problems related to the smoke detector. PHENIX had work on its Iris Scanner. STAR did trigger cabling work.
yo8-tv5 (IR8 separation bump) failed and had to be replaced.
Intermittent Quench Ground Current warnings
Spent a couple of hours trying to decouple yellow. First attempts did not work,
PHENIX is having trouble with the Blue clock. Cannot lock electronics. Channel 7 on the V124. Rob Mitchnoff and Angelika set up the clocks. Keep present store until after 10:00 AM so Waldo and Rob can look at it. Background rates in PHENIX are high. Will need Gap Cleaning, at least in Gold. Acceleration RF cavities raised from 75 to 100 kV. It improved luminosity.
Yun will work on decoupling during the evening shift today.
Present store will be kept until 11:00 AM.
STAR Beam-Beam counters like right. ZDCs look got, but the TAC is broken.
Finish rebucketing during the day using the rest of this store. Injection drift will get worked on for the next fill. And gap cleaning will get done last.
Angelika will come in to do the Blue collimators later today.
Auto-cogging not working.
Experiments will get the night shift again tonight.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email:sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |